# PoisonedRAG: Knowledge Poisoning Attacks to Retrieval-Augmented Generation of Large Language Models

### v1

@@ Abstract
Large language models (LLMs) have achieved remarkable success due to their exceptional generative capabilities. Despite their success, they also have inherent limitations such as a lack of up-to-date knowledge and hallucination. Retrieval-Augmented Generation (RAG) is a state-of-the-art technique to mitigate those limitations. In particular, given a question, RAG retrieves relevant knowledge from a knowledge database to augment the input of the LLM. For instance, the retrieved knowledge could be a set of top-$k$ texts that are most semantically similar to the given question when the knowledge database contains millions of texts collected from Wikipedia. As a result, the LLM could utilize the retrieved knowledge as the context to generate an answer for the given question. Existing studies mainly focus on improving the accuracy or efficiency of RAG, leaving its security largely unexplored. We aim to bridge the gap in this work. Particularly, we propose PoisonedRAG, a set of knowledge poisoning attacks to RAG, where an attacker could inject a few poisoned texts into the knowledge database such that the LLM generates an attacker-chosen target answer for an attacker-chosen target question. We formulate knowledge poisoning attacks as an optimization problem, whose solution is a set of poisoned texts. Depending on the background knowledge (e.g., black-box and white-box settings) of an attacker on the RAG, we propose two solutions to solve the optimization problem, respectively. Our results on multiple benchmark datasets and LLMs show our attacks could achieve 90% attack success rates when injecting 5 poisoned texts for each target question into a database with millions of texts. We also evaluate recent defenses and our results show they are insufficient to defend against our attacks, highlighting the need for new defenses.  [^1] Our code is publicly available at https://github.com/sleeepeer/PoisonedRAG
@@ 1 Introduction
Large language models (LLMs) such as GPT-3.5 [1], GPT-4 [2], and PaLM 2 [3] are widely deployed in the real world for their exceptional generative capabilities. Despite their success, they also have inherent limitations. For instance, they lack up-to-date knowledge as they are pre-trained on past data (e.g., the cutoff date for the pre-training data of GPT-4 is April 2023 [2]); they exhibit hallucination behaviors [4] (e.g., generate inaccurate content); they could have gaps of knowledge in particular domains (e.g., medical domain), especially when the data is scarce or restricted due to privacy concerns. Those limitations pose severe challenges for many real-world applications in healthcare [5, 6], finance [7], legal consulting [8, 9], scientific research [10, 11, 12], etc.
@@
Retrieval-Augmented Generation (RAG) [13, 14, 15, 16] is a state-of-the-art technique to mitigate those limitations for LLMs, which augments LLMs with external knowledge retrieved from a knowledge database. There are three components in RAG: knowledge database, retriever, and LLM. The knowledge database contains a large number of texts collected from various sources such as Wikipedia [17], financial documents [7], news articles [18], COVID-19 publications [19], to name a few. For each text in the knowledge database, the retriever uses a text encoder (e.g., BERT [20]) to compute an embedding vector for it. Given a question (e.g., “Who is the CEO of OpenAI?”) from a user, the retriever uses the text encoder to output an embedding vector for it. Then, the set of (e.g., $k$) texts (called retrieved texts) in the knowledge database whose embedding vectors have the largest similarity (e.g., cosine similarity) to that of the question are retrieved. Finally, the $k$ retrieved texts are used as the context for the LLM to generate an answer for the given question. Figure 1 shows an example of RAG.
Compared with fine-tuning [21], RAG enables LLMs to utilize external knowledge in a plug-and-play manner. Additionally, the knowledge database can be updated flexibly, e.g., it could be periodically updated to incorporate up-to-date knowledge. Because of these benefits, we have witnessed a variety of developed tools (e.g., ChatGPT Retrieval Plugin [22], LlamaIndex [23], and LangChain [24]) and real-world applications (e.g., WikiChat [25], BlueBot [26], and so on [27]) of RAG. According to 2023 Retool Report [28], more than 30% of enterprise LLM use cases now utilize the RAG technique.
@@
Existing studies [29, 30, 31, 32, 33, 34] mainly focused on improving the accuracy and efficiency of RAG. For instance, some studies [30, 31, 34] designed new retrievers such that more relevant knowledge could be retrieved for a given question. Other studies [29, 32, 33] proposed various techniques to improve the efficiency in retrieving knowledge from the knowledge database as it could contain millions of texts. However, the security of RAG is largely unexplored.
Our contribution:  In this work, we aim to bridge the gap. In particular, we propose PoisonedRAG, a set of attacks called knowledge poisoning attacks to RAG.
@@
Threat Model: In PoisonedRAG, an attacker first selects one or more questions (called target questions) and selects an arbitrary answer (called target answer) for each target question. The attacker aims to poison the knowledge database such that the LLM in a RAG generates the target answer for each target question. For instance, an attacker could mislead the LLM to generate misinformation (e.g., the target answer could be “Tim Cook” when the target question is “Who is the CEO of OpenAI?”), commercial biased answers (e.g., the answer is a particular brand over others when asked for recommendations on consumer products), and financial disinformation about markets or specific companies (e.g., falsely stating a company is facing bankruptcy when asked about its financial situation). Those attacks pose severe challenges for the deployment of RAG in many safety and reliability-critical applications such as cybersecurity, financial services, and healthcare.
Recall the three components in RAG: knowledge database, retriever, and LLM. We consider an attacker cannot access texts in the knowledge database and cannot access/query the LLM in RAG. The attacker may or may not know the retriever. With it, we consider two settings: white-box setting and black-box setting. The attacker could access the parameters of the retriever in the white-box setting (e.g., a publicly available retriever is adopted in RAG), while the attacker cannot access the parameters nor query the retriever in the black-box setting [^2] In the traditional definition of black-box attacks, the attacker could query the model. Here we consider a stronger threat model by assuming the attacker cannot even query the retriever. We still call it black-box setting for terminology consistency. . We consider the attacker could inject a few carefully crafted poisoned texts into the knowledge database. For instance, when the knowledge database contains millions of texts collected from Wikipedia, an attacker could inject poisoned texts by maliciously editing Wikipedia pages as demonstrated in the previous work [35].
@@
Overview of PoisonedRAG: We formulate crafting poisoned texts as an optimization problem. However, it is very challenging to directly solve the optimization problem (we defer details to Section 4.1). In response, we resort to heuristic solutions that involve deriving two conditions, namely retrieval condition and effectiveness condition for each poisoned text that could lead to an effective attack. The retrieval condition means a poisoned text needs to be retrieved for the target question. The effectiveness condition means the poisoned text could mislead the LLM to generate the target answer for the target question when it is used as the context. We then design attacks in both white-box and black-box settings to craft poisoned texts that simultaneously satisfy the two conditions. Our key idea is to decompose a poisoned text into two sub-texts, which are crafted to achieve two conditions, respectively. Additionally, when concatenating the two sub-texts together, they simultaneously achieve two conditions.
@@
Evaluation of PoisonedRAG: We conduct systematic evaluations of PoisonedRAG on multiple benchmark datasets (Natural Question (NQ) [36], HotpotQA [37], MS-MARCO [38]) and 8 LLMs (e.g., GPT-4 [2], LLaMA-2 [39]). We use Attack Success Rate (ASR) as the evaluation metric, which measures the fraction of target questions whose answers are attacker-desired target answers under attacks. We have the following observations from our results. First, PoisonedRAG could achieve high ASRs with a small poisoning rate. For instance, on the NQ dataset, we find that PoisonedRAG could achieve a 97% ASR by injecting 5 poisoned texts for each target question (the poisoning rate per target question is 5/2,681,468 $\approx$ 0.0002%) in the black-box setting. Second, PoisonedRAG outperforms the SOTA baselines [40, 41] and its two variants, e.g., on the NQ dataset, PoisonedRAG (black-box setting) could achieve a 97% ASR, while ASRs of baselines and two variants are less than 70%. Third, our extensive ablation studies show PoisonedRAG is robust against different hyper-parameters (in both RAG and PoisonedRAG).
@@
Defending against PoisonedRAG: We explore several defenses, including paraphrasing [42] and perplexity-based detection [42, 43, 44]. Our results show these defenses are insufficient to defend against PoisonedRAG, thus highlighting the need for new defenses. Our major contributions are as follows:
  • We propose PoisonedRAG, a set of knowledge poisoning attacks to retrieval-augmented generation of LLMs.
  • We formulate knowledge poisoning attacks as an optimization problem and design two effective solutions based on the background knowledge of an attacker.
@@
  • We conduct an extensive evaluation for PoisonedRAG on multiple benchmark datasets and LLMs. Additionally, we compare PoisonedRAG with multiple baselines.
  • We explore several defenses against PoisonedRAG. Our results show they are insufficient to defend against PoisonedRAG, highlighting the need for new defenses.
@@ 2.1 Background on RAG
Three components of a RAG system:  knowledge database, retriever, and LLM. The database contains a set of texts, e.g., collected from various sources such as Wikipedia [17], news articles [18], and financial documents [7]. For simplicity, we use $\mathcal{D}$ to denote the database that contains a set of $d$ texts, i.e., $\mathcal{D}=\{T_{1},T_{2},\cdots,T_{d}\}$, where $T_{i}$ is the $i$th text. In general, the database could be very large, e.g., it could contain millions of texts collected from Wikipedia [17]. Given a question from a user, the retriever finds the top-$k$ texts from the database that are most relevant to the given question. This set of top-$k$ texts could serve as the external knowledge to enable the LLM to generate an answer for the given question.
Two steps of a RAG system to generate an answer for a question:  Given a question $Q$, there are two steps for the LLM in the RAG system to generate an answer for it.
Step I–Knowledge Retrieval: Suppose we have two encoders in a retriever, i.e., question encoder $f_{Q}$ and text encoder $f_{T}$. The question encoder $f_{Q}$ could produce an embedding vector for an arbitrary question, while the text encoder $f_{T}$ could produce an embedding vector for each text in the knowledge database. Depending on the retriever, $f_{Q}$ and $f_{T}$ could be the same or different. Suppose we have a question $Q$, RAG first finds $k$ texts (called retrieved texts) from the database $\mathcal{D}$ that have the largest semantic similarities with the $Q$. In particular, for each $T_{i}\in\mathcal{D}$, the similarity score of $T_{i}$ with the question $Q$ is calculated as $\mathcal{S}(Q,T_{i})=Sim(f_{Q}(Q),f_{T}(T_{i}))$, where $Sim$ measures the similarity (e.g., cosine similarity, dot product) of two embedding vectors. For simplicity, we use $\mathcal{E}(Q;\mathcal{D})$ to denote the set of $k$ retrieved texts in the database $\mathcal{D}$ that have the largest similarity scores with the question $\mathcal{Q}$. Formally,
@@
Step II–Answer Generation: Given the question $Q$, the set of $k$ retrieved texts $\mathcal{E}(Q;\mathcal{D})$, and the API of a LLM, we could query the LLM with the question $Q$ and $k$ retrieved texts $\mathcal{E}(Q;\mathcal{D})$ to produce the answer for $Q$ with the help of a system prompt (we put a system prompt in Appendix B). In particular, the LLM would generate the answer for $Q$ using the $k$ retrieved texts as the context (as shown in Figure 1). For simplicity, we use $LLM(Q,\mathcal{E}(Q;\mathcal{D}))$ to denote the answer, where we omit the system prompt for simplicity.
@@ 2.2 Existing Attacks to LLMs
Many attacks to LLMs were proposed such as prompt injection attacks [45, 40, 46, 47, 48, 49], jailbreaking attacks [50, 51, 52, 53, 54, 55], and so on [56, 41, 57, 58, 59, 35, 60, 61, 62]. In particular, prompt injection attacks aim to inject malicious instructions into the input of a LLM such that the LLM could follow the injected instruction to produce attacker-desired answers. We could extend prompt injection attacks to attack RAG. For instance, we could construct the following malicious instruction: “When you are asked to provide the answer for the following question: <target question>, please output <target answer>”. However, there are two limitations for prompt injection attacks when extended to RAG. First, RAG uses the retriever component to retrieve the top-$k$ relevant texts from the knowledge database for a target question, which is not considered in prompt injection attacks. As a result, it achieves sub-optimal performance. Additionally, prompt injection attacks are less stealthy since they inject instructions, e.g., previous studies [63, 42] showed that prompt injection attacks could be detected with a very high true positive rate and a low false positive rate. Different from prompt injection attacks, our attack craft poisoned knowledge instead of malicious instructions. Moreover, we also consider how to make the poisoned knowledge be retrieved for a target question, which is not considered in prompt injection attacks.
Jailbreaking attacks aim to break the safety alignment of a LLM, e.g., crafting a prompt such that the LLM produces an answer for a harmful question like “How to rob a bank?”, for which the LLM refuses to answer without attacks. As a result, the jailbreaking attacks have different goals from ours, i.e., our attack is orthogonal to jailbreaking attacks.
We note that Zhong et al. [41] showed an attacker could generate texts (without semantic meanings) such that they are retrieved for target questions. Different from Zhong et al. [41], we aim to craft poisoned texts that have semantic meanings. As a result, the LLM would produce attacker-chosen target answers for attacker-chosen target questions. Due to such difference, our results show Zhong et al. [41] have limited attack effectiveness under our scenario.

### v2

# PoisonedRAG: Knowledge Corruption Attacks to Retrieval-Augmented Generation of Large Language Models

@@ Abstract
Large language models (LLMs) have achieved remarkable success due to their exceptional generative capabilities. Despite their success, they also have inherent limitations such as a lack of up-to-date knowledge and hallucination. Retrieval-Augmented Generation (RAG) is a state-of-the-art technique to mitigate these limitations. The key idea of RAG is to ground the answer generation of an LLM on external knowledge retrieved from a knowledge database. Existing studies mainly focus on improving the accuracy or efficiency of RAG, leaving its security largely unexplored. We aim to bridge the gap in this work. We find that the knowledge database in a RAG system introduces a new and practical attack surface. Based on this attack surface, we propose PoisonedRAG, the first knowledge corruption attack to RAG, where an attacker could inject a few malicious texts into the knowledge database of a RAG system to induce an LLM to generate an attacker-chosen target answer for an attacker-chosen target question. We formulate knowledge corruption attacks as an optimization problem, whose solution is a set of malicious texts. Depending on the background knowledge (e.g., black-box and white-box settings) of an attacker on a RAG system, we propose two solutions to solve the optimization problem, respectively. Our results show PoisonedRAG could achieve a 90% attack success rate when injecting five malicious texts for each target question into a knowledge database with millions of texts. We also evaluate several defenses and our results show they are insufficient to defend against PoisonedRAG, highlighting the need for new defenses.  [^1] Our code is publicly available at https://github.com/sleeepeer/
PoisonedRAG
@@ 1 Introduction
Large language models (LLMs) such as GPT-3.5 [1], GPT-4 [2], and PaLM 2 [3] are widely deployed in the real world for their exceptional generative capabilities. Despite their success, they also have inherent limitations. For instance, they lack up-to-date knowledge as they are pre-trained on past data (e.g., the cutoff date for the pre-training data of GPT-4 is April 2023 [2]); they exhibit hallucination behaviors [4] (e.g., generate inaccurate answers); they could have gaps of knowledge in particular domains (e.g., the medical domain). These limitations pose severe challenges for deploying LLMs in many real-world applications in healthcare [5, 6], finance [7], law [8, 9], and scientific research [10, 11, 12] fields.
@@
Retrieval-Augmented Generation (RAG) [13, 14, 15, 16] is a state-of-the-art technique to mitigate those limitations, which augments an LLM with external knowledge retrieved from a knowledge database. As shown in Figure 1, there are three components in RAG: knowledge database, retriever, and LLM. A knowledge database contains a large number of texts collected from various sources such as Wikipedia [17], financial documents [7], news articles [18], COVID-19 publications [19], to name a few. A retriever is used to retrieve a set of most relevant texts from the knowledge database for a question. With the help of a system prompt, the retrieved texts are used as the context for the LLM to generate an answer for the given question. RAG enables an LLM to utilize external knowledge in a plug-and-play manner. Moreover, RAG can reduce hallucinations and enhance the domain-specific expertise of an LLM. Due to these benefits, we have witnessed a variety of developed tools (e.g., ChatGPT Retrieval Plugin [20], LlamaIndex [21], ChatRTX [22], and LangChain [23]) and real-world applications (e.g., WikiChat [24], Bing Search [25], Clinfo.AI [26], Google Search with AI Overviews [27], Perplexity AI [28], and LLM agents [29, 30]) of RAG.
@@
Existing studies [31, 32, 33, 34, 35, 36] mainly focused on improving the accuracy and efficiency of RAG. For instance, some studies [32, 33, 36] designed new retrievers such that more relevant knowledge could be retrieved for a given question. Other studies [31, 34, 35] proposed various techniques to improve the efficiency of knowledge retrieval. However, the security of RAG is largely unexplored.
To bridge the gap, we propose PoisonedRAG, the first knowledge corruption attack to RAG.
Knowledge database as a new and practical attack surface.  In this work, we find that knowledge databases of RAG systems introduce a new and practical attack surface. In particular, an attacker can inject malicious texts into the knowledge database of a RAG system to induce an LLM to generate attacker-desired answers to user questions. For instance, when the knowledge database contains millions of texts collected from Wikipedia, an attacker could inject malicious texts by maliciously editing Wikipedia pages [37]; an attacker could also post fake news or host malicious websites to inject malicious texts when the knowledge databases are collected from the Internet; an insider can inject malicious texts into an enterprise private knowledge database.
@@
Threat model.  In PoisonedRAG, an attacker first selects one or more questions (called target questions) and selects an arbitrary answer (called target answer) for each target question. The attacker aims to inject malicious texts into the knowledge database of a RAG system such that an LLM generates the target answer for each target question. For instance, an attacker could mislead the LLM to generate misinformation (e.g., the target answer could be “Tim Cook” when the target question is “Who is the CEO of OpenAI?”), commercial biased answers (e.g., the answer is a particular brand over others when asked for recommendations on consumer products), and financial disinformation about markets or specific companies (e.g., falsely stating a company is facing bankruptcy when asked about its financial situation). These attacks pose severe challenges for deploying RAG systems in many safety and reliability-critical applications such as cybersecurity, financial services, and healthcare.
We consider an attacker cannot access texts in the knowledge database and cannot access/query the LLM in RAG. The attacker may or may not know the retriever. With it, we consider two settings: white-box setting and black-box setting. The attacker could access the parameters of the retriever in the white-box setting (e.g., a publicly available retriever is adopted in RAG), while the attacker cannot access the parameters nor query the retriever in the black-box setting. As mentioned before, we consider an attacker can inject a few malicious texts into a knowledge database of a RAG system.
@@
Overview of PoisonedRAG. We formulate crafting malicious texts as an optimization problem. However, it is very challenging to directly solve the optimization problem. In response, we resort to heuristic solutions that involve deriving two conditions, namely retrieval condition and generation condition for malicious texts that can lead to an effective attack. The retrieval condition means a malicious text can be retrieved for a target question. The generation condition means a malicious text can mislead an LLM to generate a target answer for a target question when the text is used as the context. We then design attacks in both white-box and black-box settings to craft malicious texts that simultaneously satisfy the two conditions. Our key idea is to decompose a malicious text into two sub-texts, which are crafted to achieve two conditions, respectively. Additionally, when concatenating the two sub-texts together, they simultaneously achieve these two conditions.
@@
Evaluation of PoisonedRAG.  We conduct systematic evaluations of PoisonedRAG on multiple datasets (Natural Question (NQ) [38], HotpotQA [39], MS-MARCO [40]), 8 LLMs (e.g., GPT-4 [2], LLaMA-2 [41]), and three real-world applications, including advanced RAG schemes, Wikipedia-based chatbot, and LLM agent. We use Attack Success Rate (ASR) as the evaluation metric, which measures the fraction of target questions whose answers are attacker-desired target answers under attacks. We have the following observations from our results. First, PoisonedRAG could achieve high ASRs with a small number of malicious texts. For instance, on the NQ dataset, we find that PoisonedRAG could achieve a 97% ASR by injecting 5 malicious texts for each target question into a knowledge database (with 2,681,468 clean texts) in the black-box setting. Second, PoisonedRAG outperforms the SOTA baselines [42, 43]. For instance, on the NQ dataset, PoisonedRAG (black-box setting) achieves a 97% ASR, while ASRs of 5 baselines are less than 70%. Third, our ablation studies show PoisonedRAG is robust against different hyper-parameters.
@@
Defending against PoisonedRAG.  We explore several defenses, including paraphrasing [44] and perplexity-based detection [44, 45, 46]. Our results show these defenses are insufficient to defend against PoisonedRAG, thus highlighting the need for new defenses.
Our major contributions are as follows:
  • We propose PoisonedRAG, the first knowledge corruption attack that exploit the new attack surface introduced by knowledge databases of RAG systems.
  • Our major contribution is to derive two necessary conditions for an effective attack to RAG systems. We further design PoisonedRAG to achieve these two conditions.
@@
  • We conduct an extensive evaluation for PoisonedRAG on multiple knowledge databases, retrievers, RAG schemes, and LLMs. Additionally, we compare PoisonedRAG with 5 baselines.
  • We explore several defenses against PoisonedRAG.
@@ 2.1 Background on RAG
RAG systems. There are three components for a RAG system: knowledge database, retriever, and LLM. The database contains a set of texts collected from various sources such as Wikipedia [17], news articles [18], and financial documents [7]. For simplicity, we use $\mathcal{D}$ to denote the database that contains a set of $d$ texts, i.e., $\mathcal{D}=\{T_{1},T_{2},\cdots,T_{d}\}$, where $T_{i}$ is the $i$th text. Given a question $Q$, there are two steps for the LLM in a RAG system to generate an answer for it.
Step I–Knowledge Retrieval: Suppose we have two encoders in a retriever, e.g., jointly trained question encoder $f_{Q}$ and text encoder $f_{T}$. The $f_{Q}$ produces an embedding vector for an arbitrary question, while $f_{T}$ produces an embedding vector for each text in the knowledge database. Depending on the retriever, $f_{Q}$ and $f_{T}$ could be the same or different. Suppose we have a question $Q$, RAG first finds $k$ texts (called retrieved texts) from the knowledge database $\mathcal{D}$ that are most relevant with $Q$. In particular, the similarity score of each $T_{i}\in\mathcal{D}$ with the question $Q$ is calculated as $\mathcal{S}(Q,T_{i})=Sim(f_{Q}(Q),f_{T}(T_{i}))$, where $Sim$ measures the similarity (e.g., cosine similarity, dot product) of two embedding vectors. For simplicity, we use $\mathcal{E}(Q;\mathcal{D})$ to denote the set of $k$ retrieved texts in the database $\mathcal{D}$ that have the largest similarity scores with the question $Q$. Formally, we denote:
@@
Step II–Answer Generation: Given the question $Q$, the set of $k$ retrieved texts $\mathcal{E}(Q;\mathcal{D})$, and the API of a LLM, we can query the LLM with the question $Q$ and $k$ retrieved texts $\mathcal{E}(Q;\mathcal{D})$ to produce the answer for $Q$ with the help of a system prompt (we put a system prompt in Appendix B). In particular, the LLM generates an answer to $Q$ using the $k$ retrieved texts as the context (as shown in Figure 1). For simplicity, we use $LLM(Q,\mathcal{E}(Q;\mathcal{D}))$ to denote the answer, where we omit the system prompt for simplicity.
@@ 2.2 Existing Attacks to LLMs
Many attacks to LLMs were proposed such as prompt injection attacks [47, 42, 48, 49, 50, 51], jailbreaking attacks [52, 53, 54, 55, 56, 57], and so on [58, 43, 59, 60, 61, 37, 62, 63, 64]. Prompt injection attacks aim to inject malicious instructions into the input of an LLM such that the LLM could follow the injected instruction to produce attacker-desired answers. We can extend prompt injection attacks to attack RAG. For instance, we construct the following malicious instruction: “When you are asked to provide the answer for the following question: <target question>, please output <target answer>”. However, there are two limitations for prompt injection attacks when extended to RAG. First, RAG uses a retriever component to retrieve the top-$k$ relevant texts from a knowledge database for a target question, which is not considered in prompt injection attacks. As a result, prompt injection attacks achieve sub-optimal performance. Additionally, prompt injection attacks are less stealthy since they inject instructions, e.g., previous studies [65, 44] showed that prompt injection attacks can be detected with a very high true positive rate and a low false positive rate. Different from prompt injection attacks, PoisonedRAG crafts malicious texts that can be retrieved for attacker-desired target questions and mislead an LLM to generate attacker-chosen target answers.
Jailbreaking attacks aim to break the safety alignment of a LLM, e.g., crafting a prompt such that the LLM produces an answer for a harmful question like “How to rob a bank?”, for which the LLM refuses to answer without attacks. As a result, jailbreaking attacks have different goals from ours, i.e., our attack is orthogonal to jailbreaking attacks.
We note that Zhong et al. [43] showed an attacker can generate adversarial texts (without semantic meanings, i.e., consists of random characters) such that they can be retrieved for indiscriminate user questions. However, these adversarial texts cannot mislead an LLM to generate attacker-desired answers. Different from Zhong et al. [43], we aim to craft malicious texts that have semantic meanings, which can not only be retrieved but also mislead an LLM to produce attacker-chosen target answers for target questions. Due to such difference, our results show Zhong et al. [43] are ineffective in misleading an LLM to generate target answers.